2.A quantity with an initial value of 2300 decays exponentially at a rate of 0.9% every minute. What is the value of the quantity after 441 seconds, to the nearest hundredth?


1
Expert's answer
2021-01-27T09:27:47-0500

We have the following quantities

that is to say the quantity decays exponentially

"a = 2300, \space\space r = 0.009, \space\space \space\space \space\space"

"\gamma - ? \space\space t_2 = 441s = \frac{441}{60} = 7.35 min"



"\gamma = a*(1-r)^t = 2300 * (1-0.009)^{7.35} = 2152.133"